OpenStudy (anonymous):

What is the remainder of (x3 – 6x – 9x + 3) ÷ (x - 3)

sam (.sam.):

Long division

sam (.sam.):

or f(3)=(x3 – 6x – 9x + 3)

OpenStudy (anonymous):

you have a choice you can either divide by \(x-3\) and see what you get, or you can replace \(x\) by 3 that will also give you the remainder. what .sam. said

sam (.sam.):

there's a missing x^2

OpenStudy (anonymous):

no x^3 is what it is suppose to say

sam (.sam.):

f(3)=x^3 – 6x^2 – 9x + 3??

OpenStudy (anonymous):

yes

sam (.sam.):

the 6x is it 6x^2?

OpenStudy (anonymous):

no

sam (.sam.):

then you add the like terms first f(3)=(x3 – 6x – 9x + 3) f(3)=(x3 –15x + 3)

OpenStudy (anonymous):

(x^3 – 6x – 9x + 3) ÷ (x - 3) is how it reads and the calculator gave me a strange answer

sam (.sam.):

Substitute 3 into x, f(3)=((3)^3 –15(3) + 3) f(3)=-15

sam (.sam.):

reamainder is -15

OpenStudy (callisto):

Wouldn't it be better to write f(x)=(x^3 –15x + 3)? when it's f(3) =... I think you should sub. x by 3...

OpenStudy (anonymous):

that isnt an answer option

sam (.sam.):

I think its f(x)=(x3 – 6x^2 – 9x + 3)

sam (.sam.):

f(3)=(x3 – 6x^2 – 9x + 3) remainder=-51

sam (.sam.):

is it?

OpenStudy (anonymous):

A. -51 B.-51/x - 3 C. -17/x D. -17/ x - 3

sam (.sam.):

-51

sam (.sam.):

A

sam (.sam.):

Its (x3 – 6x^2 – 9x + 3)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

 f(3) = -51